a particle of mass 2kg with velo. 10 m /s moving along a line x+y=2 its angular momentum abt pt. (2,0) will b??
1
Expert's answer
2012-12-07T03:46:41-0500
Angular momentum equals:
L=m*[r,V]
Point (2, 0) belongs to the line x + y = 2 (2 +0 = 2). This means that the
radius vector of the particle relative to the point (2, 0) is on the line x + y
= 2. The particle moves along the line x + y = 2, which means that its velocity
is directed along this line. Therefore the radius vector and velocity are
parallel, and their cross product is 0.

Answer: The angular momentum about point (2,0) will be 0.